OpenStudy (goformit100):

A coin and a die are thrown simultaneously. The probability of getting a tail and a 6 is (A)1/8 (B)1/2 (C)1/4 (D)1/12 (E)1/6

OpenStudy (anonymous):

mult the individual probs

OpenStudy (anonymous):



OpenStudy (anonymous):

p(t) = 1/2 p(6) = 1/6

OpenStudy (anonymous):

That's right pgpilot right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok so it's D :)

OpenStudy (cwrw238):

right - if events are independent the probabilities are multiplied

OpenStudy (goformit100):

ok then ?

OpenStudy (anonymous):

1/12 or D is correct

OpenStudy (anonymous):

yep

OpenStudy (goformit100):

Thank you all of you

OpenStudy (anonymous):

Yep :)